# Motives of central slope Kronecker moduli

### v2

@@ 1. Introduction
Kronecker moduli are Geometric Invariant Theory quotients parametrizing equivalence classes of tuples of linear maps up to base change. Despite being interesting in themselves as moduli spaces for a hard linear algebra problem, they have found applications in the theory of vector bundles on projective planes [7], or more generally as ambient spaces for moduli spaces of semistable sheaves on projective varieties [1], and to the tropical vertex and Gromov-Witten invariants of toric surfaces [9, 16]. Kronecker moduli can be studied with techniques of quiver moduli. For example, in the case where they are smooth and projective, their Poincaré polynomials can be computed [15]. In the special case of central slope (when the numerical parameters defining the space differ only by one) a simple formula for the Euler characteristic is proven in [20] using torus localization techniques.
The present paper starts from the second named author’s observation that this Euler characteristic coincides with a number of intervals in higher Tamari lattices [3], which makes it desirable to understand the Betti numbers of central slope Kronecker moduli better, with the ultimate aim of relating them directly to Tamari interval combinatorics. A crucial step in this direction is achieved in the present work, by describing the generating function of motives of central slope Kronecker moduli by an algebraic $q$-difference equation (in the spirit of a general result [11, 12] about algebraicity of generating series of Euler characteristic of framed Kronecker moduli). The derivation of the $q$-difference equation is made possible by utilizing various dualities of Kronecker moduli, mainly originating in the reflection functors of quiver representation theory. More precisely, our main result reads as follows:
@@
We recall all necessary notions of quiver representations, (framed) quiver moduli spaces, and their generating series of motivic invariants in Section 2. Isomorphisms of quiver moduli spaces induced by reflection functors are constructed in Section 3. Specializing these to Kronecker moduli in Section 4, we obtain crucial dualities in Corollary 4.1, Theorem 4.2. These are used in Section 5 to establish a duality of generating series of motives, which is then specialized in Section 6 to the case of central slope, and related to Tamari intervals.
Acknowledgments: The second-named author has benefited from the support of the French ANR project Charms (ANR-19-CE40-0017). The work of M. R. is supported by the DFG CRC-TRR 191 “Symplectic structures in geometry, algebra and dynamics” (281071066). The authors would like to thank the referee for suggesting several improvements of the exposition.
@@ 2.1. Basic quiver notation
acts via base change
@@ 2.2. Moduli spaces of (semi-)stable representations
where $\Delta$ is the diagonally embedded copy of $\mathbb{C}^{*}$ in $G_{\bf d}$. If non-empty, $M_{\bf d}^{\Theta\text{-}{\rm st}}(Q)$ is thus smooth and irreducible of dimension $1-\langle{\bf d},{\bf d}\rangle$. If the dimension vector ${\bf d}$ is $\Theta$-coprime, that is, $\Theta({\bf e})\not=0$ for all proper ${\bf e}\leq{\bf d}$, the stable and semistable locus coincide, thus $M_{\bf d}^{\Theta\text{-}{\rm sst}}(Q)=M_{\bf d}^{\Theta\text{-}{\rm st}}(Q)$ is smooth and projective. In this case, ${\bf d}$ is indivisible, that is, it is not a proper multiple of another dimension vector, and thus there exist universal bundles $\mathcal{V}_{i}$ of rank $d_{i}$ on $M_{\bf d}^{\Theta\text{-}{\rm sst}}(Q)$.
We make a short note on duality: changing the orientation of all arrows in the quiver $Q$ yields the opposite quiver $Q^{\rm op}$. Dualizing all vector spaces and linear maps in a quiver representation associates to $V$ a representation $V^{*}$ of $Q^{\rm op}$. On the geometric level, fixing Hermitian structures on the vector spaces $V_{i}$, we have a map $R_{\bf d}(Q)\rightarrow R_{\bf d}(Q^{\rm op})$ associating to $(f_{\alpha})_{\alpha}$ the tuple of adjoint maps $(f_{\alpha}^{*}:V_{j}\rightarrow V_{i})_{\alpha:i\rightarrow j})$, which is compatible with the $G_{\bf d}$-actions in the sense that, written symbolically,
@@ 2.3. Motives of moduli spaces
$[X]_{\rm vir}=(-\mathbb{L}^{1/2})^{-\dim X}\cdot[X]\in R.$
It will turn out that all our calculations will already take place in the subring of $R$ of rational functions in $\mathbb{L}^{1/2}$. Define the so-called motivic quantum affine space of $Q$ (see, for example, [13]) as the formal power series ring $R[[x_{i}\,|\,i\in Q_{0}]]$, with multiplication twisted by the anti-symmetrized Euler form of $Q$:
@@
In [16, Lemma 4.3] it is explained how such a factorization identity follows from a (Harder-Narasimhan type) recursive formula given in [16, Definition 4.1.(2)]. That this recursion holds on the motivic level is explained in the proof of [18, Theorem 3.5]. ∎
@@ 2.4. Framed moduli spaces
Note that the group $PG_{\widehat{\bf d}}\simeq G_{\bf d}$ is special, and $M_{{\bf d},{\bf n}}^{\Theta\text{-}{\rm fr}}(Q)$ is the geometric quotient of $R_{\widehat{\bf d}}^{\widehat{\Theta}\text{-}{\rm sst}}(\widehat{Q})$ by this group, thus
@@ 3. Reflection functors and moduli spaces
Let $i$ be a sink in $Q_{0}$. We define $s_{i}Q$ as the quiver with all arrows at $i$ reversed. Formally, we have
@@ 6. Central slope
Now we specialize to the case $k=1$ in the identities of Corollary 5.1, and note the additional identity
@@
From this equation, we will now derive an algebraic functional equation defining $F(t)$:
@@
Finally, we specialize the above functional equations to $v=1$; geometrically, this corresponds to passing from virtual motives to Euler characteristic. Our aim is to give a new proof of [20, Theorem 6.6] avoiding the iterated torus fixed point localization techniques used there.
@@
$\frac{m-1}{d((m-2)d+1)}{{(m-1)^{2}d+m-2}\choose{d-1}}.$
Using [3], we have thus rederived:
@@
[20] The Euler characteristic of $K_{d,d-1}^{(m)}$ equals the number of length $d$ intervals in the $(m-2)$-Tamari lattice.

### v3

@@ 1. Introduction
Kronecker moduli are Geometric Invariant Theory quotients parametrizing equivalence classes of tuples of linear maps up to change of basis. Despite being interesting in themselves as moduli spaces for a hard linear algebra problem, they have found applications in the theory of vector bundles on projective planes [7], or more generally as ambient spaces for moduli spaces of semistable sheaves on projective varieties [1], and to the tropical vertex and Gromov-Witten invariants of toric surfaces [9, 15]. Kronecker moduli can be studied with techniques of quiver moduli. For example, in the case where they are smooth and projective, their Poincaré polynomials can be computed [14]. In the special case of central slope (when the numerical parameters defining the space differ only by one) a simple formula for the Euler characteristic is proven in [19] using torus localization techniques.
The present paper starts from the second named author’s observation that this Euler characteristic coincides with a number of intervals in higher Tamari lattices [3], which makes it desirable to understand the Betti numbers of central slope Kronecker moduli better, with the ultimate aim of relating them directly to Tamari interval combinatorics. A crucial step in this direction is achieved in the present work, by describing the generating function of motives of central slope Kronecker moduli by an algebraic $q$-difference equation (in the spirit of a general result [11] about algebraicity of generating series of Euler characteristic of framed Kronecker moduli). The derivation of the $q$-difference equation is made possible by utilizing various dualities of Kronecker moduli, mainly originating in the reflection functors of quiver representation theory. More precisely, our main result reads as follows:
@@
We recall all necessary notions of quiver representations, (framed) quiver moduli spaces, and their generating series of motivic invariants in Section 2. Isomorphisms of quiver moduli spaces induced by reflection functors are constructed in Section 3. Specializing these to generalized Kronecker quivers, and thus to Kronecker moduli, in Section 4, we obtain crucial dualities in Corollary 4.1, Theorem 4.2. These are used in Section 5 to establish a duality of generating series of motives, which is then specialized in Section 6 to the case of central slope, and related to Tamari intervals. Moreover, Section 6 discusses several other formulas for the motives and their generating series, as well as examples.
Acknowledgments: The second-named author has benefited from the support of the French ANR project Charms (ANR-19-CE40-0017). The work of M. R. is supported by the DFG CRC-TRR 191 “Symplectic structures in geometry, algebra and dynamics” (281071066). The authors would like to thank the referees for suggesting several improvements of the exposition.
@@ 2.1. Basic quiver notation
acts via change of basis
@@ 2.2. Moduli spaces of (semi-)stable representations
where $\Delta$ is the diagonally embedded copy of $\mathbb{C}^{*}$ in $G_{\bf d}$. If non-empty, $M_{\bf d}^{\Theta\text{-}{\rm st}}(Q)$ is thus smooth and irreducible of dimension $1-\langle{\bf d},{\bf d}\rangle$. If the dimension vector ${\bf d}$ is $\Theta$-coprime, that is, $\Theta({\bf e})\not=0$ for all proper ${\bf e}\leq{\bf d}$ (componentwise inequality), the stable and semistable locus coincide, thus $M_{\bf d}^{\Theta\text{-}{\rm sst}}(Q)=M_{\bf d}^{\Theta\text{-}{\rm st}}(Q)$ is smooth and projective. In this case, ${\bf d}$ is indivisible, that is, it is not a proper multiple of another dimension vector, and thus there exist universal bundles $\mathcal{V}_{i}$ of rank $d_{i}$ on $M_{\bf d}^{\Theta\text{-}{\rm sst}}(Q)$.
Changing the orientation of all arrows in the quiver $Q$ yields the opposite quiver $Q^{\rm op}$. Dualizing all vector spaces and linear maps in a quiver representation associates to $V$ a representation $V^{*}$ of $Q^{\rm op}$. On the geometric level, fixing ($\mathbb{C}$-bilinear) non-degenerate symmetric bilinear forms on the vector spaces $V_{i}$, we have a map $R_{\bf d}(Q)\rightarrow R_{\bf d}(Q^{\rm op})$ associating to $(f_{\alpha})_{\alpha}$ the tuple of adjoint maps $(f_{\alpha}^{*}:V_{j}\rightarrow V_{i})_{\alpha:i\rightarrow j})$, which is compatible with the $G_{\bf d}$-actions in the sense that, written symbolically,
@@ 2.3. Motives of moduli spaces
$[X]_{\rm vir}=(-\mathbb{L}^{1/2})^{-\dim X}\cdot[X]\in R$
(in the present context, working with virtual motives has the advantage of making several formulas more symmetric). It will turn out that all our calculations will already take place in the subring of $R$ generated by the rational functions in $\mathbb{L}^{1/2}$. Define the so-called motivic quantum affine space of $Q$ (see, for example, [12]) as the formal power series ring $R[[x_{i}\,|\,i\in Q_{0}]]$, with multiplication twisted by the anti-symmetrized Euler form of $Q$:
@@
In [15, Lemma 4.3] it is explained how such a factorization identity follows from a (Harder-Narasimhan type) recursive formula given in [15, Definition 4.1.(2)]. That this recursion holds on the motivic level is explained in the proof of [17, Theorem 3.5]. ∎
@@ 2.4. Framed moduli spaces
Note that the group $PG_{\widehat{\bf d}}\simeq G_{\bf d}$ is special (see [4, Definition 3.7]), and $M_{{\bf d},{\bf n}}^{\Theta\text{-}{\rm fr}}(Q)$ is the geometric quotient of $R_{\widehat{\bf d}}^{\widehat{\Theta}\text{-}{\rm sst}}(\widehat{Q})$ by this group, thus
@@ 3. Reflection functors and moduli spaces
We continue to denote by $Q$ a finite acyclic quiver. Let $i$ be a sink in $Q_{0}$. We define $s_{i}Q$ as the quiver with all arrows at $i$ reversed. Formally, we have
@@ 6. Central slope
Now we specialize to the case $k=1$ in the identities of Corollary 5.1; note that this corresponds to the case of dimension vectors of slope one (which we call the central slope), thus $d=e$. We note the additional identity
@@
Taking the definition of $F(t)$ as a generating series of virtual motives and comparing coefficients in the previous formula, we find the following recursion which allows for practical calculations:
Corollary 6.3.
The virtual motives $m_{d}=[K_{d,d}^{(m),{\rm fr}}]_{\rm vir}$ are given recursively by $m_{0}=1$ and
$m_{d}=\frac{[\mathbb{P}^{(m-1)d}]_{\rm vir}}{[\mathbb{P}^{d-1}]_{\rm vir}}\sum_{d_{1}+\ldots+d_{m-1}=d-1}v^{\sum_{i}(m-2i)d_{i}}\prod_{i}m_{d_{i}}.$
From Corollary 6.2, we will now derive an algebraic functional equation defining $F(t)$:
@@
Finally, we specialize the above functional equations to $v=1$; geometrically, this corresponds to passing from virtual motives to Euler characteristic. Our aim is to give a new proof of [19, Theorem 6.6] avoiding the iterated torus fixed point localization techniques used there.
@@
$\frac{m-1}{d((m-2)d+1)}{{(m-1)^{2}d+m-2}\choose{d-1}},$
which rederives [19, Theorem 6.6].
Using the calculation of the number of intervals in generalized Tamari lattices in [3], we have thus proved:
@@
The Euler characteristic of $K_{d,d-1}^{(m)}$ equals the number of intervals in the $(m-2)$-Tamari lattice of index $d$.